# biblio.bib

### v1

@@ 5.2 Test on a Traffic Circle
plus the reversed arcs (Figure 4).
$x$$y$$z_{1}$$z_{2}$$z_{3}$$z_{4}$$z_{5}$$z_{6}$$z_{7}$$z_{8}$$\gamma_{1}$$\gamma_{2}$$\gamma_{3}$$\gamma_{4}$$\gamma_{5}$$\gamma_{6}$$\gamma_{7}$$\gamma_{8}$$\gamma_{9}$$\gamma_{10}$$\gamma_{11}$$\gamma_{12}$
@@ 5.2.1 Hamiltonian Dependent on $s$
Since the exact critical value for this problem is not explicitly known, we take $\hat{c}=2.59\cdot 10^{-1}$ as the reference critical value, computed by running Algorithm 2 with parameters $\Delta x=1.25\cdot 10^{-2}$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and $\varepsilon=\Delta x/10$, and rounding to the first three significant figures.
We carry out the study of the algorithms, starting with Algorithm 1. The algorithm is tested with $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and with $2000$ fixed iterations to be executed. The corresponding results are presented in Table 7 and Figure 5. We compare these results with those in columns 8–10 of Table 8, obtained by implementing the algorithm with the same time step $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ but setting the tolerance $\varepsilon=\Delta x/10$. The observed errors are approximately the same. As illustrated in Figure 5, the sequences $(\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k})/2$ decrease in $k$ and the errors $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-\hat{c}\right|$ reach a plateau, whose value decreases in $\Delta x$. This confirms that choosing $\varepsilon=\Delta x/10$ is not restrictive, as already analyzed in Section 5.1.1.
We point out that, while Figure 5 (right) shows decreasing errors in $k$ toward a plateau, Figure 5 (center) reports a different behavior. For instance, in the test with $\Delta x=2.50\cdot 10^{-2}$, the error reaches the lowest point at iteration 165 before converging and stabilizing at a certain value. In this test, when $k=165$, the distance between $\overline{c}^{\Delta}_{k}$ and $\underline{c}^{\Delta}_{k}$ is not smaller than $2\varepsilon$, as can be seen in Figure 5 (left and right), however their average is very close to the reference critical value.
We continue our study setting $\varepsilon=\Delta x/10$, $\Delta t=\Delta x^{5/6}$, $\Delta t=\Delta x/2$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and comparing the results given by Algorithm 1 and Algorithm 2, shown in Table 8 and Table 9, respectively. What stands out is that Algorithm 2 requires fewer iterations than Algorithm 1. In fact, Algorithm 2 yields an average reduction in the number of iterations as following: $67\%$ for $\Delta t=\Delta x^{5/6}$ (column 2 of Table 8 and Table 9), $84\%$ for $\Delta t=\Delta x/2$ (column 5 of Table 8 and Table 9) and $87\%$ for $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ (column 8 of Table 8 and Table 9).
Figure 5: Algorithm 1. Test on the traffic circle with Hamiltonian dependent on $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and $2000$ fixed iterations. The third graph refers to the test with $\Delta x=2.50\cdot 10^{-2}$.
${\Delta x}$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$
$2.00\cdot 10^{-1}$ $2000$ $0.41$ $1.53\cdot 10^{-1}$
@@
Table 7: Algorithm 1. Test on the traffic circle with Hamiltonian dependent on $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and $2000$ fixed iterations.
${\Delta x}$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$
@@
Table 8: Algorithm 1. Test on the traffic circle with Hamiltonian dependent on $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$.
${\Delta x}$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$
@@
Table 9: Algorithm 2. Test on the traffic circle with Hamiltonian dependent on $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$.
@@ 5.2.2 Hamiltonian Independent of $s$
For this problem we can choose $\beta_{0}=7.5$ and, as in the previous case, the exact critical value is unknown. We set the reference critical value $\hat{c}=-1.50$ as the value computed by Algorithm 2 with $\Delta x=1.25\cdot 10^{-2}$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$, $\varepsilon=\Delta x/10$ and rounded to the first three significant figures.
We test Algorithm 1 with $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ and $2000$ fixed iterations to be executed (Table 10 and Figure 6). The current problem is independent of $s$ and, as detailed in Section 5.1.2, the approximation error introduced by the semi-Lagrangian scheme is smaller than the estimate in Theorem 4.1. As a result, the errors $\left|\frac{\overline{c}^{\Delta}_{k}+\underline{c}^{\Delta}_{k}}{2}-\hat{c}\right|$ decrease and do not reach a plateau, as shown in Figure 6. Since $(\overline{c}^{\Delta}_{k}-\underline{c}^{\Delta}_{k})/2$ decreases in $k$ (see Figure 6), when we implement Algorithm 1 with $\varepsilon=\Delta x/10$ (columns 8–10 of Table 11), the decreasing trend of the errors as $\Delta x$ decreases is a consequence of the chosen tolerance.
We conduct some tests with $\varepsilon=\Delta x/10$ and compare the outcomes of Algorithm 1 and Algorithm 2, which are presented in Table 11 and Table 12, respectively. The tests are performed with three different values of $\Delta t$, given by $\Delta t=\Delta x^{5/6}$, $\Delta t=\Delta x/2$ and $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$. Numerical simulations show that, by implementing Algorithm 2 instead of Algorithm 1, we achieve an average reduction in the number of iterations by: $66\%$ for $\Delta t=\Delta x^{5/6}$ (column 2 of Table 11 and Table 12), $83\%$ for $\Delta t=\Delta x/2$ (column 5 of Table 11 and Table 12) and $88\%$ for $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ (column 8 of Table 11 and Table 12).
Figure 6: Algorithm 1. Test on the traffic circle with Hamiltonian independent of $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ and $2000$ fixed iterations.
${\Delta x}$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$
$2.00\cdot 10^{-1}$ $2000$ $-1.50$ $4.20\cdot 10^{-4}$
$1.00\cdot 10^{-1}$ $2000$ $-1.50$ $3.44\cdot 10^{-4}$
$5.00\cdot 10^{-2}$ $2000$ $-1.50$ $3.10\cdot 10^{-4}$
$2.50\cdot 10^{-2}$ $2000$ $-1.50$ $2.77\cdot 10^{-4}$
Table 10: Algorithm 1. Test on the traffic circle with Hamiltonian independent of $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ and $2000$ fixed iterations.
${\Delta x}$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$
$2.00\cdot 10^{-1}$ $244$ $-1.51$ $1.23\cdot 10^{-2}$ $232$ $-1.50$ $2.97\cdot 10^{-3}$ $232$ $-1.50$ $3.62\cdot 10^{-3}$
$1.00\cdot 10^{-1}$ $479$ $-1.51$ $5.73\cdot 10^{-3}$ $465$ $-1.50$ $1.26\cdot 10^{-3}$ $465$ $-1.50$ $1.48\cdot 10^{-3}$
$5.00\cdot 10^{-2}$ $952$ $-1.50$ $2.84\cdot 10^{-3}$ $938$ $-1.50$ $5.94\cdot 10^{-4}$ $938$ $-1.50$ $6.61\cdot 10^{-4}$
$2.50\cdot 10^{-2}$ $1894$ $-1.50$ $9.82\cdot 10^{-4}$ $1878$ $-1.50$ $2.75\cdot 10^{-4}$ $1878$ $-1.50$ $2.95\cdot 10^{-4}$
$1.25\cdot 10^{-2}$ $3774$ $-1.50$ $4.29\cdot 10^{-4}$ $3756$ $-1.50$ $1.30\cdot 10^{-4}$ $3756$ $-1.50$ $1.36\cdot 10^{-4}$
Table 11: Algorithm 1. Test on the traffic circle with Hamiltonian independent of $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$.
${\Delta x}$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$ $k$ $c^{\Delta}$ $|c^{\Delta}-\hat{c}|$
$2.00\cdot 10^{-1}$ $76$ $-1.51$ $9.31\cdot 10^{-3}$ $40$ $-1.50$ $7.33\cdot 10^{-4}$ $29$ $-1.50$ $6.66\cdot 10^{-4}$
$1.00\cdot 10^{-1}$ $126$ $-1.50$ $4.36\cdot 10^{-3}$ $77$ $-1.50$ $2.19\cdot 10^{-4}$ $55$ $-1.50$ $3.20\cdot 10^{-5}$
$5.00\cdot 10^{-2}$ $271$ $-1.50$ $2.06\cdot 10^{-3}$ $153$ $-1.50$ $4.24\cdot 10^{-5}$ $109$ $-1.50$ $8.55\cdot 10^{-6}$
$2.50\cdot 10^{-2}$ $717$ $-1.50$ $6.91\cdot 10^{-4}$ $306$ $-1.50$ $1.21\cdot 10^{-5}$ $218$ $-1.50$ $1.19\cdot 10^{-5}$
$1.25\cdot 10^{-2}$ $1656$ $-1.50$ $2.18\cdot 10^{-4}$ $616$ $-1.50$ $2.94\cdot 10^{-6}$ $437$ $-1.50$ $5.15\cdot 10^{-7}$
Table 12: Algorithm 2. Test on the traffic circle with Hamiltonian independent of $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$.
@@
The purpose of this Appendix is to prove Theorem 3.11, extending [Pozza23, Theorem 5.2] to the case where condition (D) does not hold. We point out that our proof does not require condition (H2), nor the flux limiter to be minimal, but only that
$c_{x}\leq c,\qquad\text{for every }x\in\mathbf{V}.$ (21)
To prove it we need some preliminary results, adapted from [Pozza23] to our case. Through this Appendix we assume (21) to be in force.
@@
[…]
Proof.
The proof is the same of [Pozza23, Proposition 5.7], even if the Aubry sets are different from our setting. ∎
[…]
$u_{1}(x)=(\mathcal{S}(t)u_{1})(x)+ct\leq(\mathcal{S}(t)\phi)(x)+ct\leq u_{2}(x),\qquad\text{for any }(x,t)\in\Gamma\times\mathds{R}^{+}.$ (22)
Thanks to this and the uniform continuity of $\mathcal{S}(t)\phi$, see Theorem 3.6, the Arzelà–Ascoli Theorem yields that given any positive diverging sequence $\{t_{n}\}$ it is always possible to extract a subsequence $\{t_{n_{k}}\}$ so that $\mathcal{S}(t_{n_{k}})\phi+ct_{n_{k}}$ converges uniformly to some $f\in C(\Gamma)$. We denote with $\omega_{\mathcal{S}}(\phi)$ the set made up by the uniform limits of $\mathcal{S}(t)\phi+ct$. We further set the semilimits
@@
where the supremum and the infimum are taken over the sequences $\{x_{n}\}_{n\in\mathds{N}}$ converging to $x$ and the positive diverging sequences $\{t_{n}\}_{n\in\mathds{N}}$. In view of the uniform continuity of $\mathcal{S}(t)\phi$, both $\underline{\phi}$ and $\overline{\phi}$ are continuous and the sequences $\{x_{n}\}$ may be chosen identically equal to $x$. It follows that
$\underline{\phi}(x)=\sup\{f(x):f\in\omega_{\mathcal{S}}(\phi)\}\quad\text{and}\quad\overline{\phi}(x)=\inf\{f(x):f\in\omega_{\mathcal{S}}(\phi)\}.$ (25)

### v2

@@ 5.2 Test on a Traffic Circle
plus the reversed arcs (Figure˜4).
[…]
@@ 5.2.1 Hamiltonian Dependent on $s$
Since the exact critical value for this problem is not explicitly known, we take $\hat{c}=2.59\cdot 10^{-1}$ as the reference critical value, computed by running Algorithm˜2 with parameters $\Delta x=1.25\cdot 10^{-2}$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and $\varepsilon=\Delta x/10$, and rounding to the first three significant figures.
[…]
Figure 5: Algorithm˜1. Test on the traffic circle with Hamiltonian dependent on $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and $2000$ fixed iterations. The third graph refers to the test with $\Delta x=2.50\cdot 10^{-2}$.
[…]
$\cdot 10^{-1} 200020$$0$ $0.41$ $1.53\cdot 10^{-1}$
@@
Table 7: Algorithm˜1. Test on the traffic circle with Hamiltonian dependent on $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and $2000$ fixed iterations.
[…]
@@
Table 8: Algorithm˜1. Test on the traffic circle with Hamiltonian dependent on $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$.
We point out that, while Figure˜5 (right) shows decreasing errors in $k$ toward a plateau, Figure˜5 (center) reports a different behavior. For instance, in the test with $\Delta x=2.50\cdot 10^{-2}$, the error reaches the lowest point at iteration 165 before converging and stabilizing at a certain value. In this test, when $k=165$, the distance between $\overline{c}^{\Delta}_{k}$ and $\underline{c}^{\Delta}_{k}$ is not smaller than $2\varepsilon$, as can be seen in Figure˜5 (left and right), however their average is very close to the reference critical value.
We continue our study setting $\varepsilon=\Delta x/10$, $\Delta t=\Delta x^{5/6}$, $\Delta t=\Delta x/2$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ and comparing the results given by Algorithm˜1 and Algorithm˜2, shown in Table˜8 and Table˜9, respectively. What stands out is that Algorithm˜2 requires fewer iterations than Algorithm˜1. In fact, Algorithm˜2 yields an average reduction in the number of iterations as following: $67\%$ for $\Delta t=\Delta x^{5/6}$ (column 2 of Table˜8 and Table˜9), $84\%$ for $\Delta t=\Delta x/2$ (column 5 of Table˜8 and Table˜9) and $87\%$ for $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$ (column 8 of Table˜8 and Table˜9).
[…]
@@
Table 9: Algorithm˜2. Test on the traffic circle with Hamiltonian dependent on $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/9.5$.
@@ 5.2.2 Hamiltonian Independent of $s$
For this problem we can choose $\beta_{0}=7.5$ and, as in the previous case, the exact critical value is unknown. We set the reference critical value $\hat{c}=-1.50$ as the value computed by Algorithm˜2 with $\Delta x=1.25\cdot 10^{-2}$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$, $\varepsilon=\Delta x/10$ and rounded to the first three significant figures.
@@
Figure 6: Algorithm˜1. Test on the traffic circle with Hamiltonian independent of $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ and $2000$ fixed iterations.
[…]
$\cdot 10^{-1} 200020$$0$ 1.50-1.50 $4.20\cdot 10^{-4}$
$1.00\cdot 10^{-1}$ $2000$ 1.50-1.50 $3.44\cdot 10^{-4}$
$5.00\cdot 10^{-2}$ $2000$ 1.50-1.50 $3.10\cdot 10^{-4}$
$2.50\cdot 10^{-2}$ $2000$ 1.50-1.50 $2.77\cdot 10^{-4}$
Table 10: Algorithm˜1. Test on the traffic circle with Hamiltonian independent of $s$, $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ and $2000$ fixed iterations.
We conduct some tests with $\varepsilon=\Delta x/10$ and compare the outcomes of Algorithm˜1 and Algorithm˜2, which are presented in Table˜11 and Table˜12, respectively. The tests are performed with three different values of $\Delta t$, given by $\Delta t=\Delta x^{5/6}$, $\Delta t=\Delta x/2$ and $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$. Numerical simulations show that, by implementing Algorithm˜2 instead of Algorithm˜1, we achieve an average reduction in the number of iterations by: $66\%$ for $\Delta t=\Delta x^{5/6}$ (column 2 of Table˜11 and Table˜12), $83\%$ for $\Delta t=\Delta x/2$ (column 5 of Table˜11 and Table˜12) and $88\%$ for $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$ (column 8 of Table˜11 and Table˜12).
[…]
$2.00\cdot 10^{-1}$ $244$ 1.51-1.51 $1.23\cdot 10^{-2}$ $232$ 1.50-1.50 $2.97\cdot 10^{-3}$ $232$ 1.50-1.50 $3.62\cdot 10^{-3}$
$1.00\cdot 10^{-1}$ $479$ 1.51-1.51 $5.73\cdot 10^{-3}$ $465$ 1.50-1.50 $1.26\cdot 10^{-3}$ $465$ 1.50-1.50 $1.48\cdot 10^{-3}$
$5.00\cdot 10^{-2}$ $952$ 1.50-1.50 $2.84\cdot 10^{-3}$ $938$ 1.50-1.50 $5.94\cdot 10^{-4}$ $938$ 1.50-1.50 $6.61\cdot 10^{-4}$
$2.50\cdot 10^{-2}$ $1894$ 1.50-1.50 $9.82\cdot 10^{-4}$ $1878$ 1.50-1.50 $2.75\cdot 10^{-4}$ $1878$ 1.50-1.50 $2.95\cdot 10^{-4}$
$1.25\cdot 10^{-2}$ $3774$ 1.50-1.50 $4.29\cdot 10^{-4}$ $3756$ 1.50-1.50 $1.30\cdot 10^{-4}$ $3756$ 1.50-1.50 $1.36\cdot 10^{-4}$
Table 11: Algorithm˜1. Test on the traffic circle with Hamiltonian independent of $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$.
[…]
$2.00\cdot 10^{-1}$ $76$ 1.51-1.51 $9.31\cdot 10^{-3}$ $40$ 1.50-1.50 $7.33\cdot 10^{-4}$ $29$ 1.50-1.50 $6.66\cdot 10^{-4}$
$1.00\cdot 10^{-1}$ $126$ 1.50-1.50 $4.36\cdot 10^{-3}$ $77$ 1.50-1.50 $2.19\cdot 10^{-4}$ $55$ 1.50-1.50 $3.20\cdot 10^{-5}$
$5.00\cdot 10^{-2}$ $271$ 1.50-1.50 $2.06\cdot 10^{-3}$ $153$ 1.50-1.50 $4.24\cdot 10^{-5}$ $109$ 1.50-1.50 $8.55\cdot 10^{-6}$
$2.50\cdot 10^{-2}$ $717$ 1.50-1.50 $6.91\cdot 10^{-4}$ $306$ 1.50-1.50 $1.21\cdot 10^{-5}$ $218$ 1.50-1.50 $1.19\cdot 10^{-5}$
$1.25\cdot 10^{-2}$ $1656$ 1.50-1.50 $2.18\cdot 10^{-4}$ $616$ 1.50-1.50 $2.94\cdot 10^{-6}$ $437$ 1.50-1.50 $5.15\cdot 10^{-7}$
Table 12: Algorithm˜2. Test on the traffic circle with Hamiltonian independent of $s$ and $\varepsilon=\Delta x/10$. Columns 2–4 refer to $\Delta t=\Delta x^{5/6}$, columns 5–7 refer to $\Delta t=\Delta x/2$ and columns 8–10 refer to $\Delta t=(\min_{\gamma\in\mathcal{E}^{+}}\Delta_{\gamma}x)/7.5$.
Appendix A Proof of Theorem 3.7
˜3.7 has been proved in [24, Theorem 5.2] under the additional assumption
  1. (D)
    for any $\gamma\in\mathcal{E}$ with $a_{\gamma}=c=a_{0}$ the map $s\mapsto\min\limits_{p\in\mathds{R}}H_{\gamma}(s,p)$ is constant in $0,|\gamma|]$.
This assumption relies on knowing the value of $c$ in advance, which is clearly too restrictive for our purposes. Under this condition the extended Aubry set $\widetilde{\mathcal{A}}_{\Gamma}$ consists of the support of a collection of arcs, while in our case it could even be made up of a countable number of disjointed pieces of arcs or just a single point. However, the proof of [24, Theorem 5.2] can be readily extended to our case using the same methods, with mostly straightforward modifications. Thus, we will just sketch the proof of Theorem˜3.7, highlighting the main steps.
First, the same arguments used in the proof of [24, Proposition 5.7] prove the following result.
@@
Let $w$ be a subsolution to ($\mathcal{H}$Jcc), then $(x,t)\mapsto w(x)-ct$ is a subsolution to ($\mathcal{H}$JE). If $w$ is also a solution to ($\mathcal{H}$Jcc) in $\Gamma\setminus\left(\widetilde{\mathcal{A}}_{\Gamma}\setminus\mathcal{A}_{\Gamma}\right)$, then $w-ct=\mathcal{S}(t)w$ on $\Gamma\times\mathds{R}^{+}$.
Setting $u$ as in ˜6 and
we get, according to the comparison principle [27, Theorem 7.1], Propositions˜A.1, 3.2 and 3.5, that
$w(x)$ $\leq(\mathcal{S}(t)w)(x)+ct\leq u(x),$ $\text{for any $(x,t)\in\Gamma\times\mathds{R}^{+}$},$ (20)
$w(x)$ $=(\mathcal{S}(t)w)(x)+ct=u(x),$ $\text{for any $(x,t)\in\widetilde{\mathcal{A}}_{\Gamma}\times\mathds{R}^{+}$}.$
Let us define the semilimits
@@
where the supremum and the infimum are taken over the sequences $\{x_{n}\}_{n\in\mathds{N}}$ converging to $x$ and the positive diverging sequences $\{t_{n}\}_{n\in\mathds{N}}$. Arguing as in [24, Proposition 5.13] we get
@@
Given $\phi\in C(\Gamma)$, let $\underline{\phi}$ and $\overline{\phi}$ be as defined in (21) and (22), respectively. Then $\underline{\phi}$ and $\overline{\phi}$ are a subsolution to ($\mathcal{H}$Jcc) on $\Gamma$ and a supersolution to ($\mathcal{H}$Jcc) on $\Gamma\setminus\widetilde{\mathcal{A}}_{\Gamma}$, respectively.
Defining the semilimits $\underline{w}$ and $\overline{w}$ as in ˜21 and 22 with $\phi$ replaced by $w$, respectively, we get from ˜20
$\overline{w}(x)$ $\leq\underline{w}(x)\leq u(x),\qquad\text{for any $x\in\Gamma$},$
$\overline{w}(x)$ $=\underline{w}(x)=u(x),\qquad\text{for any $x\in\widetilde{\mathcal{A}}_{\Gamma}$},$
then Proposition˜A.2 and the comparison principle for the eikonal equation [23, Theorem 5.3] yield that $\overline{w}=\underline{w}=u$ on $\Gamma$, which in turn shows
$\lim_{t\to\infty}(\mathcal{S}(t)w)(x)+ct=u(x),\qquad\text{for every $x\in\Gamma$}.$ (23)
Next we have by Proposition˜3.2 that